# spacing=nonfrench

### 04 Mar 2026 v1

@@ Abstract.
To a height-$1$ formal group defined over the ring of integers of a finite extension $K$ of $\mathbb{Q}_{p}$ is attached its $p$-adic Tate module, which is a crystalline character of $\operatorname{Gal}(\overline{K}/K)$ of Hodge-Tate weight $1$. This association is an equivalence. We prove, over extensions whose ramification index is relatively prime to $p^{2}-p$, that the set of consistent sequences attached to a height-$1$ commuting pair $(f,u)$ of noninvertible and invertible formal power series is a crystalline character of weight $1$, for which $f$ is an endomorphism. As a result, we deduce a proof of a conjecture of Lubin in new cases.
@@ Introduction
Pairs of analytic transformations of a $p$-adic open unit disk with a fixed point at 0 that commute under composition were studied by Lubin in [Lub94] under the name of $p$-adic dynamical systems. He pointed out in ibidem that “for an invertible series to commute with a noninvertible series, there must be a formal group somehow in the background”. This observation has initiated a wave of results in this direction, see for instance [Li96], [Li02], [Sar05], [Sar10], [SS13], [Ber19]. More recently, $p$-adic Hodge theory has proved to be an efficient way for tackling this problem, as in [Ber16], [Spe18], [Poy24], and this paper is written in that vein.
@@
Throughout this article, we refer to the above statement as Lubin’s conjecture. The height of the dynamical system together with its ring of definition is a common way to decompose this conjecture into cases. The height-$1$ case over $\mathbb{Z}_{p}$ is solved by Specter in [Spe18] where he recovered the underlying Lubin-Tate power series. Shortly after, Berger solved in [Ber19] all the cases of finite height over $\mathbb{Z}_{p}$ using only $p$-adic analysis. In this article, we solve the height-$1$ cases over extensions whose ramification index is relatively prime to $p^{2}-p$, provided that the linear coefficient of the invertible series lies in $\mathbb{Z}_{p}$. The remaining cases of Lubin’s conjecture are still open nowadays. Our approach consists in generalizing Specter’s analysis of the set of consistent sequences of the dynamical system which yields a Galois character, recovering the missing $\mathbb{Z}_{p}$-module structure on it, and, from this data, reconstructing the latent formal group via integral $p$-adic Hodge-theoretic methods.
@@
Let $p$ be any prime number and fix an algebraic closure of $\mathbb{Q}_{p}$ containing all our extensions. Define $\mathbb{C}_{p}$ to be its $p$-adic completion, with open unit disk $\mathfrak{m}_{\mathbb{C}_{p}}\!$. If $E$ is a finite extension of $\mathbb{Q}_{p}$, we let $\mathcal{O}_{E}$ be its ring of integers, with maximal ideal $\mathfrak{m}_{E}$, and we let $e_{E}$ be the ramification index of the extension, or $e$ if it is unambiguous. Given a ring $R$, we define $\mathscr{S}_{0}(R)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=XR[\![X]\!]$ to be the set of formal power series over $R$ without constant term. Composition induces a law on $\mathscr{S}_{0}(R)$ making it into a noncommutative monoid, and a necessary and sufficient condition on $s\in\mathscr{S}_{0}(R)$ to be invertible is that $s^{\prime}(0)$ is a unit of $R$. We deal exclusively with rings $R$ arising from a $p$-adic situation, that is, $p$-adic fields, their integers and finite fields of characteristic $p$.
@@
Before stating our main theorem, we remind the reader of some elementary facts about formal groups, see also [Haz12] for an extended presentation. By a formal group, we mean a one-dimensional commutative formal group law. Given a formal group $F$ defined over the ring of integers of a finite extension $K/\mathbb{Q}_{p}$, its ring of $\mathcal{O}_{K}$-endomorphisms, namely $\operatorname{End}_{\mathcal{O}_{K}}(F)$, is a composition-commuting family of noninvertible and invertible formal power series, provided with an embedding of $\mathbb{Z}_{p}$ sending $m\in\mathbb{Z}_{p}$ to $m]_{F}$, the multiplication-by-$m$ map in $F$. The map taking the linear coefficient in $\operatorname{End}_{\mathcal{O}_{K}}(F)$ is injective. If we further assume that $F$ is of height $1$ and $K$ is a nontrivial extension of $\mathbb{Q}_{p}$, then $F$ is not Lubin-Tate and this is the major difference with Specter’s case. What remains true is that $\operatorname{End}_{\mathcal{O}_{K}}(F)$ is a free $\mathbb{Z}_{p}$-algebra of rank $1$, implying that the linear coefficient of every $\mathcal{O}_{K}$-endomorphism of $F$ lies in $\mathbb{Z}_{p}$. We will refer to this fact as (1). An important data attached to the formal group $F$ is its $p$-adic Tate module
@@
where $\Lambda_{n}(F)$ is the group of $p^{n}$-torsion points of $F$ in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ for every integer $n\geqslant 0$. This is a crystalline character of the absolute Galois group $G_{K}$ of $K$ of Hodge-Tate weight $1$, meaning that $\operatorname{T}_{p}(F)$ is a free $\mathbb{Z}_{p}$-module of rank $1$ provided with a compatible action of $G_{K}$ that is crystalline of weight $1$. According to [SW13, Corollary 6.2.3], what makes it relevant is that the functor
@@
is an equivalence, given that those of weight $1$ correspond to connected $p$-divisible groups, that is, formal groups [Tat67, Proposition 1]. Sufficiently explicit functors to recover a formal group from an object on the right-hand side can be extracted from the work of Kisin [Kis06], [Kis09] and Zink [Zin01], [Zin02] in integral $p$-adic Hodge theory.
This equivalence plays a key role in this article. Motivated by the above discussion, and given a power series $f\in\mathscr{S}_{0}(\mathcal{O}_{K})$ that has exactly $p$ roots in the open unit disk of $\mathbb{C}_{p}$, whose iterates have only simple roots and such that $f^{\prime}(0)$ is a uniformizer in $\mathbb{Z}_{p}$, we consider the $G_{K}$-set
@@
of $f$-consistent sequences of elements in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ (see [Lub94, p. 329]), where $\Lambda_{n}(f)$ is now the set of roots of $f^{\circ n}$ in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ for all integers $n\geqslant 0$. If Lubin’s conjecture holds, the latter should be the $p$-adic Tate module of an integral formal group once $f$ commutes with a nontorsion invertible $u\in\mathscr{S}_{0}(\mathcal{O}_{K})$. Therefore, there should exist a compatible $\mathbb{Z}_{p}$-module structure on $T_{f}$ turning it into a crystalline character of weight $1$ and admitting $f$ as an endomorphism. This is what we achieve to recover, for a restriction of $G_{K}$ from $K$ to a finite Galois extension $L$, when $(e_{K},p^{2}-p)=1$. Based on (1), we assume as well that $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}$. Although this assumption is reasonable, the author regrets not knowing if it can be deduced from the original hypotheses. The $\mathbb{Z}_{p}[G_{L}]$-module structure on $T_{f}$ allows to recover a formal group over $\mathcal{O}_{L}$ for which $(f,u)$ is a pair of endomorphisms, and the uniqueness of the latent formal group immediately implies the descent to $\mathcal{O}_{K}$.
Main theorem (Theorem 3.1).
Let $K/\mathbb{Q}_{p}$ be a finite extension satisfying $(e_{K},p^{2}-p)=1$. Let $(f,u)$ be a commuting pair of elements in $\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $f$ has exactly $p$ roots in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ and all roots of its iterates are simple, with $f^{\prime}(0)$ a uniformizer in $\mathbb{Z}_{p}\!$ and $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}$ nontorsion. Then there exists a formal group $F$ over $\mathcal{O}_{K}$ such that $f,u\in\operatorname{End}_{\mathcal{O}_{K}}(F)$.
@@
Our work has a nontrivial intersection with the one of Berger in [Ber19]. If we assume $(f,u)$ to be a commuting pair over $\mathcal{O}_{K}$ as in our main theorem with $(e,p^{2}-p)=1$ and that e<p1e<p-1, then the same arguments as for [Ber19, Corollary A] can be repeated to show that $f$ and $u$ are endomorphisms of a formal group defined over $\mathcal{O}_{K}$. We do not claim any responsibility in this result since it is just an observation.
@@
In the first section, we recall some invariants and relations attendant to a $p$-adic dynamical system $(f,u)$ from [Lub94]. Using the main theorem of [Li96], we determine the decreasing part of the Newton polygon of any iterate of $f$. This yields information on the roots of the iterates of $f$ and also on the periodic points of $u$. The situation slightly differs from [Spe18] but a similar path can be followed modulo some new adjustments and arguments. The second section is devoted to the recovery of the Tate module structure on the set $T_{f}$ attached to $(f,u)$. In §2.1, we mimic as in [Spe18] the Galois action on $T_{f}$ by means of the action of the $\mathbb{Z}_{p}$-iterates of $u$, in order to attach a character $\chi_{f}\colon G_{L}\rightarrow\mathbb{Z}_{p}^{\times}$ to the pair $(f,u)$, where $L$ is the finite extension obtained by adjoining to $K$ the fixed points of $u$. In §2.2 and §2.3, we expose the material used in §2.4 to construct a crystalline period of $\chi_{f}$. We show that it has weight $1$. This property implies the existence of a formal group $H$ over $\mathcal{O}_{L}$ satisfying $\operatorname{T}_{p}(H)\simeq\mathbb{Z}_{p}(\chi_{f})$. Although $H$ need not be latent to our dynamical system, it should be isomorphic to it. Therefore, we rely on $H$ as a cornerstone of our constructions. In §2.5, we begin by identifying the counterpart of $(f,u)$ in $H$. This allows us, in a second time, to perform a transport of structure from $\operatorname{T}_{p}(H)$ to $T_{f}$ and turn $T_{f}$ into a $\mathbb{Z}_{p}[G_{L}]$-module. In the third section, we move backward until we get a formal group. From [Kis06], we consider the associated Breuil-Kisin module in §3.1, extend it into a window in §3.2 as in [Kis09], and transform it into a connected $p$-divisible group in §3.3 as in [Zin01]. At every stage, we track the behavior of $f$ on the new object, allowing us to conclude in §3.4 that the formal group associated to the connected $p$-divisible group is the latent one.
@@ Acknowledgments
I am sincerely grateful to my advisor, Prof. Maja Volkov, for her guidance and invaluable advice, to Giovanni Bosco for helpful conversations on $p$-adic Hodge theory, and to Prof. Laurent Berger and Prof. Jonathan Lubin for their support and help.
@@ 1. The dynamical system
Let $K$ be a finite extension of $\mathbb{Q}_{p}$ satisfying $(e_{K},p^{2}-p)=1$. Throughout this article, we assume $(f,u)$ to be a commuting pair of formal power series in $\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $f$ has exactly $p$ roots in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ and all roots of its iterates are simple, with $f^{\prime}(0)$ a uniformizer in $\mathbb{Z}_{p}$ and $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}$ nontorsion. Up to replacing $u$ by one of its iterates, we may assume that $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. In this section, we review the immediate data at hand in such a dynamical system which will play a crucial role in the subsequent work.
@@ 1.1. Logarithm
In his seminal article [Lub94], Lubin showed that in the presence of a formal power series $s\in\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $s^{\prime}(0)$ is nonzero nor a root of unity, there is a unique invertible power series
@@
satisfying $\operatorname{Log}_{s}^{\prime}(0)=1$ and $\operatorname{Log}_{s}(s(X))=s^{\prime}(0)\operatorname{Log}_{s}(X)$, called the logarithm of $s$. Every power series satisfying the same condition as above and commuting with $s$ has the same logarithm. Denote by $\operatorname{Log}_{f}$ the logarithm of $(f,u)$; it converges on $\mathfrak{m}_{\mathbb{C}_{p}}\!$. If $f$ and $u$ are endomorphisms of an integral formal group, then $\operatorname{Log}_{f}$ is the logarithm of this formal group and it is well known that its derivative has coefficients in $\mathcal{O}_{K}$. It turns out that this is true even for $p$-adic dynamical systems [Ber19, Lemma 2.2]. The conjugation
$F(X,Y)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Log}_{f}^{\circ-1}(\operatorname{Log}_{f}(X)+\operatorname{Log}_{f}(Y))$
is the unique formal group for which $f$, $u$ are endomorphisms, but it is defined a priori over $K$ rather than $\mathcal{O}_{K}$. Therefore, solving Lubin’s conjecture amounts to show that $F$ is defined over $\mathcal{O}_{K}$. For every $m\in\mathbb{Z}_{p}$, we let
$m]_{f}(X)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Log}_{f}^{\circ-1}(m\operatorname{Log}_{f}(X))\in\mathscr{S}_{0}(K)$
be the unique power series over $K$ commuting with $f$ and acting on the tangent space at 0 by $m$, which is nothing but the multiplication-by-$m$ map in $F$.

### 22 Apr 2026 v2

@@ Abstract.
Let $K/\mathbb{Q}_{p}$ be a finite extension whose ramification index is coprime to $p^{2}-p$. We study height-one commuting pairs $(f,u)$ of noninvertible and invertible formal power series defined over the ring of integers $\mathcal{O}_{K}$ of $K$. We begin by extracting a crystalline character of weight $1$ from the $\operatorname{Gal}(\overline{K}/K)$-set $T_{f}$ of $f$-consistent sequences. This character is used in order to equip $T_{f}$ with a $\mathbb{Z}_{p}$-module structure for which $f$ is an endomorphism. We then apply explicit functors in integral $p$-adic Hodge theory to $T_{f}$ to recover a formal group defined over $\mathcal{O}_{K}$ for which $(f,u)$ is a pair of endomorphisms. This proves new cases of a conjecture of Lubin.
Key words and phrases:
$p$-adic dynamical system, $p$-adic Hodge theory, formal group
@@ Introduction
Pairs of analytic transformations of a $p$-adic open unit disk with a fixed point at 0 that commute under composition were studied by Lubin in [Lub94] under the name of $p$-adic dynamical systems. He pointed out in ibidem that “for an invertible series to commute with a noninvertible series, there must be a formal group somehow in the background”. This observation has initiated a wave of results in this direction, see for instance [Li96], [Li02], [Sar05], [Sar10], [SS13], [Ber19]. More recently, $p$-adic Hodge theory has proved to be an efficient way for tackling this problem, as in the work of [Ber16], [Spe18], [Poy24], and this paper is written in this vein.
@@
Throughout this article, we refer to the above statement as Lubin’s conjecture. The height of the dynamical system together with its ring of definition is a common way to decompose this conjecture into cases. The case of height $1$ over $\mathbb{Z}_{p}$ is solved by Specter in [Spe18] where he recovered the underlying Lubin-Tate power series. Shortly after, Berger solved in [Ber19] all the cases of finite height over $\mathbb{Z}_{p}$ using only $p$-adic analysis. In this article, we solve the height-one cases over cyclo-tame extensions, that is, extensions whose ramification index is coprime to $p^{2}-p$, provided that the linear coefficient of the invertible series lies in $\mathbb{Z}_{p}$. The remaining cases of Lubin’s conjecture are still open nowadays. Our approach consists in generalizing Specter’s analysis in [Spe18] of the set of consistent sequences of the dynamical system which yields a Galois character, recovering the missing $\mathbb{Z}_{p}$-module structure on it and, from this data, retrieving the latent formal group of the dynamical system via integral $p$-adic Hodge-theoretic methods.
@@
Let $p$ be a prime number and fix an algebraic closure of $\mathbb{Q}_{p}$ containing all our extensions. Define $\mathbb{C}_{p}$ to be its $p$-adic completion, with open unit disk $\mathfrak{m}_{\mathbb{C}_{p}\!}$. When $E$ is a finite extension of $\mathbb{Q}_{p}$, we let $\mathcal{O}_{E}$ be its ring of integers, with maximal ideal $\mathfrak{m}_{E}$, and we let $e_{E}$ be the ramification index of the extension or simply $e$ if it is unambiguous. Given a ring $R$, we define $\mathscr{S}_{0}(R)=XR[\![X]\!]$ to be the set of formal power series over $R$ without constant term. Composition induces a law on $\mathscr{S}_{0}(R)$ making it into a noncommutative monoid, and a necessary and sufficient condition on $s\in\mathscr{S}_{0}(R)$ to be invertible is that $s^{\prime}(0)$ is a unit of $R$. We deal exclusively with rings $R$ arising from a $p$-adic situation, that is, $p$-adic fields, their integers and finite fields of characteristic $p$.
@@
Before stating our main theorem, we remind the reader of some elementary results about formal groups, see also [Haz12] for an extended presentation. By a formal group, we mean a one-dimensional commutative formal group law. Given a formal group $F$ defined over the ring of integers $\mathcal{O}_{K}$ of a finite extension $K$ of $\mathbb{Q}_{p}$, its ring of $\mathcal{O}_{K}$-endomorphisms, namely $\operatorname{End}_{\mathcal{O}_{K}}(F)$, is a composition-commuting family of noninvertible and invertible formal power series, provided with an embedding of $\mathbb{Z}_{p}$ sending $m\in\mathbb{Z}_{p}$ to $m]_{F}$, the multiplication-by-$m$ endomorphism in $F$. The map taking the linear coefficient in $\operatorname{End}_{\mathcal{O}_{K}}(F)$ is injective. If we further assume that $F$ has height $1$ and $K$ is a nontrivial extension of $\mathbb{Q}_{p}$, then $F$ is not a Lubin-Tate formal group and this is the major difference with Specter’s case. What remains true is that $\operatorname{End}_{\mathcal{O}_{K}}(F)$ is a free $\mathbb{Z}_{p}$-algebra of rank $1$, implying that the linear coefficient of every $\mathcal{O}_{K}$-endomorphism of $F$ lies in $\mathbb{Z}_{p}$. We will refer to this fact as (1). An important data attached to the formal group $F$ is its $p$-adic Tate module
@@
where $F[p^{n}]$ is the group of $p^{n}$-torsion points of $F$ in $\mathfrak{m}_{\mathbb{C}_{p}\!}$ for every integer $n\geqslant 0$. This is a crystalline character of the absolute Galois group $G_{K}$ of $K$ of Hodge-Tate weight $1$, meaning that $\operatorname{T}_{p}(F)$ is a free $\mathbb{Z}_{p}$-module of rank $1$ equipped with an action of $G_{K}$ that is crystalline of weight $1$. By [SW13, Corollary 6.2.3] what makes this object relevant is that the functor
@@
is an equivalence. Indeed, by [Tat67] height-one formal groups correspond to height-one connected $p$-divisible groups, giving rise to crystalline characters of weight $1$, and [Bre00, Theorem 5.3.2] when $p\neq 2$ and [Kis06, Corollary 2.2.6] for any $p$ show that any such character arises from such a $p$-divisible group. Now sufficiently explicit functors to recover a formal group from the right-hand side can be extracted from the work of Kisin [Kis06], [Kis09] and Zink [Zin01], [Zin02] in integral $p$-adic Hodge theory.
Let $f\in\mathscr{S}_{0}(\mathcal{O}_{K})$ be a power series whose reduction modulo $\mathfrak{m}_{K}$ is nonzero, such that the roots of $f$ and all of its iterates in $\mathfrak{m}_{\mathbb{C}_{p}\!}$ are simple, $f$ has exactly $p$ of them and with $f^{\prime}(0)$ a uniformizer in $\mathbb{Z}_{p}$. Motivated by the above discussion, we consider the $G_{K}$-set
@@
of $f$-consistent sequences of elements in $\mathfrak{m}_{\mathbb{C}_{p}\!}$ (see [Lub94, p. 329]), where $\Lambda_{n}(f)$ is the set of roots of $f^{\circ n}$ in $\mathfrak{m}_{\mathbb{C}_{p}\!}$ for every integer $n\geqslant 0$. If Lubin’s conjecture holds, $T_{f}$ should be the $p$-adic Tate module of an integral formal group once $f$ commutes with a nontorsion invertible power series $u\in\mathscr{S}_{0}(\mathcal{O}_{K})$. Therefore there should exist a compatible $\mathbb{Z}_{p}$-module structure on $T_{f}$ turning it into a crystalline character of weight $1$ and admitting $f$ as an endomorphism. This is what we achieve to recover, for a restriction of $G_{K}$ from $K$ to a finite extension $L$, when $(e_{K},p^{2}-p)=1$. Based on (1), we assume as well that $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}$. Although this assumption is reasonable, the author regrets not knowing if it can be deduced from the original hypotheses. The $\mathbb{Z}_{p}[G_{L}]$-module structure on $T_{f}$ allows to recover a formal group over $\mathcal{O}_{L}$ for which $(f,u)$ is a pair of endomorphisms. A uniqueness argument then implies that this formal group is defined over $\mathcal{O}_{K}$.
Main theorem (3.1).
Let $K$ be a finite extension of $\mathbb{Q}_{p}\!$ satisfying $(e_{K},p^{2}-p)=1$. Let $(f,u)$ be a commuting pair of elements in $\mathscr{S}_{0}(\mathcal{O}_{K})\!$ such that $f\bmod\mathfrak{m}_{K}\neq 0$, the roots of $f$ and all of its iterates in $\mathfrak{m}_{\mathbb{C}_{p}\!}$ are simple and $f$ has $p$ of them, with $f^{\prime}(0)$ a uniformizer in $\mathbb{Z}_{p}\!$ and $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}\!$ nontorsion. Then there exists a formal group $F$ defined over $\mathcal{O}_{K}\!$ such that $f,u\in\operatorname{End}_{\mathcal{O}_{K}}(F)$.
@@
Our work has a nontrivial intersection with the one of Berger in [Ber19]. If we assume $(f,u)$ to be a commuting pair over $\mathcal{O}_{K}$ as in our main theorem with $(e,p^{2}-p)=1$ and that e<p1e<p-1, then the same arguments as for [Ber19, Corollary A] can be repeated to show that $(f,u)$ is a pair of endomorphisms of a formal group defined over $\mathcal{O}_{K}$.
@@
In the first section, we begin by recalling some invariants and relations attendant to a $p$-adic dynamical system $(f,u)$ from [Lub94]. Using the main theorem of [Li96], we determine the decreasing part of the Newton polygon of any iterate of $f$. This yields information on the roots of the iterates of $f$ and also on the periodic points of $u$. The situation slightly differs from the case of [Spe18] but a similar path can be followed modulo some new adjustments and arguments. The second section is devoted to the recovery of the $p$-adic Tate module structure on the set $T_{f}$ attached to $(f,u)$. In §2.1, we mimic as in [Spe18] the Galois action on $T_{f}$ by means of the action of the $\mathbb{Z}_{p}$-iterates of $u$ in order to attach a character $\chi_{f}\colon G_{L}\rightarrow\mathbb{Z}_{p}^{\times}$ to $(f,u)$, where $L$ is the finite extension obtained by adjoining to $K$ the fixed points of $u$. In §2.2 and §2.3, we expose the material used in §2.4 in order to construct a crystalline period of $\chi_{f}$. We then prove that it has weight $1$. These properties carry the existence of a formal group $H$ over $\mathcal{O}_{L}$ whose Tate module is isomorphic to $\mathbb{Z}_{p}(\chi_{f})$. Although $H$ need not be latent to the dynamical system, it should be isomorphic to it. Thus we rely on $H$ as a cornerstone of our constructions. In §2.5, we begin by identifying the counterpart of $(f,u)$ in $H$. This allows us, in a second time, to transport the structure of $\operatorname{T}_{p}(H)$ to $T_{f}$ and to turn the latter into a $\mathbb{Z}_{p}[G_{L}]$-module for which $f$ is an endomorphism. In the third section, we move backward in explicit integral $p$-adic Hodge theory until we get a formal group. From [Kis06] we consider the associated Breuil-Kisin module in §3.1, then extend it into a window as in [Kis09] in §3.2, and transform it into a connected $p$-divisible group as in [Zin01] in §3.3. At every stage, we track the behavior of $f$ on the new object, allowing us to conclude in §3.4 that the formal group associated to the connected $p$-divisible group is the latent one.
@@ Acknowledgments
I am sincerely grateful to my advisor Maja Volkov for her guidance and invaluable advice and to Giovanni Bosco for helpful conversations on $p$-adic Hodge theory. I am also particularly grateful to Laurent Berger for many valuable comments on an earlier version of this paper and to Jonathan Lubin for his kind support.
@@ 1. The dynamical system
Let $K/\mathbb{Q}_{p}$ be a finite extension satisfying $(e_{K},p^{2}-p)=1$. Throughout this article, we assume $(f,u)$ to be a commuting pair of formal power series in $\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $f\bmod\mathfrak{m}_{K}\neq 0$, the roots of $f$ and all of its iterates in $\mathfrak{m}_{\mathbb{C}_{p}\!}$ are simple and $f$ has $p$ of them, where $f^{\prime}(0)$ is a uniformizer in $\mathbb{Z}_{p}$ and $u^{\prime}(0)\in\mathbb{Z}_{p}^{\times}$ is nontorsion. Up to replacing $u$ by one of its $\mathbb{Z}$-iterates, we may assume that $u^{\prime}(0)\in 1+p\mathbb{Z}_{p}$. In this section, we review the immediate data at hand in such a dynamical system which will play a crucial role in the subsequent work.
@@ 1.1. Logarithm
In his seminal article [Lub94], Lubin showed that, in the presence of a formal power series $s\in\mathscr{S}_{0}(\mathcal{O}_{K})$ such that $s^{\prime}(0)$ is nonzero nor a root of unity, there exists a unique invertible power series
@@
satisfying $\operatorname{Log}_{s}^{\prime}(0)=1$ and $\operatorname{Log}_{s}(s(X))=s^{\prime}(0)\operatorname{Log}_{s}(X)$, called the logarithm of $s$. Every power series satisfying the same condition as above and commuting with $s$ has the same logarithm. Denote by $\operatorname{Log}_{f}$ the logarithm of the commuting pair $(f,u)$; it converges on $\mathfrak{m}_{\mathbb{C}_{p}\!}$. If $f$, $u$ are endomorphisms of an integral formal group, then $\operatorname{Log}_{f}$ is the logarithm of this formal group and it is well known that its derivative has coefficients in $\mathcal{O}_{K}$. It turns out that this is true even for our $p$-adic dynamical system $(f,u)$ by [Ber19, Lemma 2.2]. The conjugated
$F(X,Y)=\operatorname{Log}_{f}^{\circ-1}(\operatorname{Log}_{f}(X)+\operatorname{Log}_{f}(Y))$
is the unique formal group for which $f$ and $u$ are endomorphisms, but it is defined a priori over $K$ rather than $\mathcal{O}_{K}$. Therefore solving Lubin’s conjecture amounts to show that $F$ is defined over $\mathcal{O}_{K}$. For every $m\in\mathbb{Z}_{p}$, we let
$m]_{f}(X)=\operatorname{Log}_{f}^{\circ-1}(m\operatorname{Log}_{f}(X))\in\mathscr{S}_{0}(K)$
be the unique series over $K$ commuting with $f$ and acting on the tangent space at 0 by $m$, which is nothing but the multiplication-by-$m$ endomorphism in $F$.